A movie theater is 3 blocks due north of a supermarket and a beauty parlor is 4 blocks due east of the movie theater. How many blocks long is the street that runs directly from the supermarket to the beauty parlor?
Answer by mananth(16949)

they form a rught triangle.
parlor to movie theater acts a hypotenuse
3^2+4^2=hyp^2
9+16 = hyp^2
25= hyp^2
5 = hyp
...
parlor to movie theater is 5 blocks distance
...
m.ananth@hotmail.ca